Type of relations for aRb if and only if a<= b+1


1
Expert's answer
2021-05-07T09:47:50-0400

Solution:

Consider relation R defined on integers (Z).

Reflexive:

"aRa\Rightarrow a\le a+1\Rightarrow True"

Thus, reflexive.

Symmetric:

"aRb\Rightarrow a\le b+1
\\ bRa \Rightarrow b\le a+1\Rightarrow False"

Thus, not symmetric.

Transitive:

"aRb\Rightarrow a\le b+1
\\bRc\Rightarrow b\le c+1"

"\Rightarrow a\le (c+1)+1=c+2\ or\ a\le c+1\Rightarrow aRc"

Thus, transitive.

Hence, given relation R is reflexive and transitive only for all "a,b,c\in Z" .